Session: Oct 2012 (session 3)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Yebes (Daniel Valero at Sun Oct 28 15:34:50 2012) Success:
Not kown problems

Medicina (Roda at Sun Oct 28 15:45:38 2012) Success:

Torun (K. Borkowski at Mon Oct 29 06:30:18 2012) Success:
No known problems.

Westerbork (Geert Kuper at Mon Oct 29 09:30:14 2012) Success:
missed first 35 minutes due to backend problems. (IVC RT8) Started at 13:05 UT.

Onsala (Michael Lindqvist at Mon Oct 29 12:14:51 2012) Success:
No known problems.

Effelsberg (Uwe Bach at Mon Oct 29 12:15:44 2012) Failure:RECE
Failure of the secondary focus receiver cooling just after n12c4. All receivers including 6cm got warm. Repair started right away on Sunday, but didn't finish within eg066f observing time.

Jodrell_Bank (Alastair Gunn at Mon Nov 12 15:48:32 2012) Success:
No known problems.